### v1

@@ 3.2 A Sequence-Form Characterization of Normal-Form Perfect Equilibria
The profile $\gamma^{*}(\varepsilon)\in\Lambda(\varepsilon)$ is a Nash equilibrium of $\Gamma_{s}(\varepsilon)$ if and only if, for each player $i\in N$ and $j\in M_{i},a\in A(I^{j}_{i})$, it holds that $\gamma^{*i}(\varepsilon;w^{i}_{I^{j}_{i}}a)=\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)$ whenever $w^{i}_{I^{j}_{i}}a$ fails to be a best-response sequence to $\gamma^{*}(\varepsilon)$.
@@
For a given $\gamma^{i}\in\Lambda^{i}$ of player $i\in N$, let $y(\gamma^{i},\varepsilon)=(y(\gamma^{i},\varepsilon;w^{i}):w^{i}\in W^{i})$, where $y(\gamma^{i},\varepsilon;w^{i})=\eta^{i}(\varepsilon;w^{i})+(1-\eta^{i}(\emptyset))\gamma^{i}(w^{i})$. It follows that $y(\cdot,\varepsilon)$ is a bijection from $\Lambda^{i}$ to $\Lambda^{i}(\varepsilon)$. Consequently, we obtain the following equivalent form of the optimization problem (5),
$\begin{array}[]{rl}\max\limits_{\gamma^{i}}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}(\varepsilon))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)-\gamma^{i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &0\leq\gamma^{i}(w^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (7)
@@
For player $i\in N$, there exists some $\gamma^{*i}\in\Lambda^{i}$ such that $y(\gamma^{*i},\varepsilon)=\gamma^{*i}(\varepsilon)$ and solves against $\gamma^{*-i}(\varepsilon)$ the optimization problem (7). By combining the above discussion with Lemma 3, we conclude that for any $j\in M_{i},a\in A(I^{j}_{i})$ where $w^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{*}(\varepsilon)$, it holds that $\gamma^{*i}(w^{i}_{I^{j}_{i}}a)=0$. Therefore, $\gamma^{*i}(\varepsilon;w^{i}_{I^{j}_{i}}a)=y(\gamma^{*i},\varepsilon;w^{i}_{I^{j}_{i}}a)=\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)$. This completes the proof. ∎
@@
Assume there exists a sequence of perturbed games $\{\Gamma_{s}(\varepsilon^{k})\}_{k=1}^{\infty}$ with $\gamma^{*}(\varepsilon^{k})$ being a Nash equilibrium for each $\Gamma_{s}(\varepsilon^{k})$ and $\lim_{k\to\infty}\gamma^{*}(\varepsilon^{k})=\gamma^{*}$. Let $\{\sigma^{k}\}_{k=1}^{\infty}$ be a sequence of totally mixed strategies with $(\sigma^{k},\gamma^{*}(\varepsilon^{k}))\in T$ and $\lim_{k\to\infty}\sigma^{k}=\sigma^{*}$. Note that $\sigma^{k}$ does not necessarily equal $\sigma(\gamma^{*}(\varepsilon^{k}))$. For any pure strategy $s^{i}$ of player $i\in N$, if $u^{i}(s^{i},\sigma^{k-i})<u^{i}(\tilde{s}^{i},\sigma^{k-i})$ holds for some $\tilde{s}^{i}\in S^{i}$, then by Lemma 3, there exists a sequence $w^{i}_{I^{j}_{i}}a$ for $j\in M_{i},a\in A(I^{j}_{i})$ such that $s^{i}(w^{i}_{I^{j}_{i}}a)=1$ and $w^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{*}(\varepsilon^{k})$. Accordingly, Lemma 4 implies that $\gamma^{*i}(\varepsilon^{k};w^{i}_{I^{j}_{i}}a)=\sum_{s^{i}\in S^{i}}s^{i}(w^{i}_{I^{j}_{i}}a)\sigma^{ki}(s^{i})=\eta^{i}(\varepsilon^{k};w^{i}_{I^{j}_{i}}a)$,which leads to $\sigma^{ki}(s^{i})\leq\eta^{i}(\varepsilon^{k};w^{i}_{I^{j}_{i}}a)\leq\varepsilon^{k}$. Therefore the sufficiency follows immediately from Definition 1.
@@
Conversely, assume that $\sigma^{*}$ is a normal-form perfect equilibrium of $\Gamma$ and $\gamma^{*}=\gamma(\sigma^{*})$. Then, there exists a sequence of totally mixed strategies $\{\sigma(\varepsilon^{k})\}_{k=1}^{\infty}$ such that $\lim_{k\to\infty}\varepsilon^{k}=0$ and $\lim_{k\to\infty}\sigma(\varepsilon^{k})=\sigma^{*}$, where each $\sigma(\varepsilon^{k})$ is an $\varepsilon^{k}$-normal form perfect equilibrium. Consider a specific $\sigma(\varepsilon^{k})$ with sufficiently large $k$, if $u^{i}(s^{i},\sigma^{-i}(\varepsilon^{k}))<u^{i}(\tilde{s}^{i},\sigma^{-i}(\varepsilon^{k}))$ holds for some $\tilde{s}^{i}\in S^{i}$, then $\sigma^{i}(\varepsilon^{k};s^{i})\leq\varepsilon^{k}$. Let $\gamma^{k}=\gamma(\sigma(\varepsilon^{k}))$ and $\tilde{\varepsilon}^{k}=\max_{i\in N}|S^{i}|\varepsilon^{k}$, we construct a perturbed game $\Gamma_{s}(\tilde{\varepsilon}^{k})$ with $\gamma^{k}$ being a Nash equilibrium. The perturbation $\eta(\tilde{\varepsilon}^{k})=(\eta^{i}(\tilde{\varepsilon}^{k};w^{i}):i\in N,w^{i}\in W^{i})$ adheres to the system (4), defined for $(j,a)\in D_{i}$ as follows,
$\eta^{i}(\tilde{\varepsilon}^{k};w^{i}_{I^{j}_{i}}a)=\left\{\begin{array}[]{ll}\gamma^{ki}(w^{i}_{I^{j}_{i}}a)&\text{if $w^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{k}$},\\ \varepsilon^{k}&\text{otherwise.}\end{array}\right.$
@@
It can be observed that $0\leq\eta^{i}(\tilde{\varepsilon}^{k};w^{i}_{I^{j}_{i}}a)\leq\tilde{\varepsilon}^{k}$. Furthermore, the recursive expressions of $\eta(\tilde{\varepsilon}^{k})$ and $\gamma^{k}$ ensure that $\eta^{i}(\tilde{\varepsilon}^{k};w^{i}_{I^{j}_{i}}a)=\gamma^{ki}(w^{i}_{I^{j}_{i}}a)$ holds for all $i\in N,j\in M_{i}$,and $a\in A(I^{j}_{i})$, provided that $w^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{k}$. Applying Lemma 4, we conclude that $\gamma^{k}$ is a Nash equilibrium for $\Gamma_{s}(\tilde{\varepsilon}^{k})$. As $\lim_{k\to\infty}\gamma^{k}=\gamma^{*}$, the proof is complete. ∎
@@ 4.1 An Artificial Game and Equilibrium Convergence Analysis
Let $\varepsilon=1/\max_{i\in N,j\in M_{i}}|A(I^{j}_{i})|$ and $\eta^{0}=(\eta^{0i}(w^{i}):i\in N,w^{i}\in W^{i})$ be a given vector that satisfies the system (4). Furthermore, let $\gamma^{0}=(\gamma^{0i}(w^{i}):i\in N,w^{i}\in W^{i})$ denote a given realization plan profile with $\gamma^{0i}(w^{i})\geq\eta^{0i}(w^{i})$, which serves as a starting point for the smooth path discussed later. The proposed method also requires the following functions, defined over the interval $[0,2]$,
@@
$\begin{array}[]{l}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ 0\leq\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;j\in M_{i},a\in A(I^{j}_{i}),\end{array}$ (8)
@@
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-c(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}(t)))\\ &+c(t)\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j}_{i}}a))\\ &+\theta(t)\sum\limits_{(j,a)\notin D_{i}}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;w^{i}_{I^{j}_{i}}a))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (9)
@@
$\begin{array}[]{l}(1-c(t))g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))+\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ (\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j}_{i}}a))\lambda^{i}(w^{i}_{I^{j}_{i}}a)=c(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a),\\ \hskip 105.2751pt\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j}_{i}}a)<\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i},\\ \gamma^{i}(t;w^{i}_{I^{j}_{i}}a)\lambda^{i}(w^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a),\;0<\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\notin D_{i},\end{array}$ (10)
where $\zeta^{i}_{I^{j}_{i}}(a)=\sum_{{j_{q}}\in M_{i}(w^{i}_{I^{j}_{i}}a)}\nu^{i}_{I^{j_{q}}_{i}}$. It can be observed that $\gamma^{*}(t)$ solves the optimization problem (9) against itself if and only if there exists a unique pair $(\lambda^{*},\nu^{*})$ such that, along with $\gamma^{*}(t)$, they fulfill the system (10). For values of $t\in(0,1]$, the condition $\gamma^{*}(t)\in\text{int}(\Lambda)$ ensures that $\sigma(\gamma^{*}(t))$ is well-defined and $(\sigma(\gamma^{*}(t)),\gamma^{*}(t))\in T$.
@@
Let $V$ denote the set of $v\in\mathbb{R}^{n_{0}}$ satisfying $f_{1}(v)\leq 0,\cdots,f_{l}(v)\leq 0,p_{1}(v)=0,\cdots,p_{q}(v)=0$, where each $f_{i}$ and $p_{j}$ is a polynomial with real coefficients. Suppose that $V$ is nonempty. Then there exist constants $\tau>0,\kappa>0$ and $\kappa^{\prime}>0$ such that $dist(v,V)\leq\tau(1+\|v\|)^{\kappa^{\prime}}(\|[f(v)]_{+}\|+\|p(v)\|)^{\kappa}$ for any $v\in\mathbb{R}^{n_{0}}$. Here $dist(\cdot,\cdot)$ denotes the Euclidean distance function between two sets, $f(v)=(f_{1}(v),\cdots,f_{l}(v))^{\top}$, $p(v)=(p_{1}(v),\cdots,p_{q}(v))^{\top}$, and $[\cdot]_{+}$ denotes the positive part of a vector.

### v3

@@ 3.2 A Sequence-Form Characterization of Normal-Form Perfect Equilibria
The profile $\gamma^{*}(\varepsilon)\in\Lambda(\varepsilon)$ is a Nash equilibrium of $\Gamma_{s}(\varepsilon)$ if and only if, for each player $i\in N$ and $j\in M_{i},a\in A(I^{j}_{i})$, it holds that $\gamma^{*i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)=\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)$ whenever $\varpi^{i}_{I^{j}_{i}}a$ fails to be a best-response sequence to $\gamma^{*}(\varepsilon)$.
@@
For a given $\gamma^{i}\in\Lambda^{i}$ of player $i\in N$, let $y(\gamma^{i},\varepsilon)=(y(\gamma^{i},\varepsilon;\varpi^{i}):\varpi^{i}\in W^{i})$, where $y(\gamma^{i},\varepsilon;\varpi^{i})=\eta^{i}(\varepsilon;\varpi^{i})+(1-\eta^{i}(\emptyset))\gamma^{i}(\varpi^{i})$. It follows that $y(\cdot,\varepsilon)$ is a bijection from $\Lambda^{i}$ to $\Lambda^{i}(\varepsilon)$. We establish in the following an optimization problem (7),
$\begin{array}[]{rl}\max\limits_{\gamma^{i}}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}(\varepsilon))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &0\leq\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (7)
@@
For player $i\in N$, there exists some $\gamma^{*i}\in\Lambda^{i}$ such that $y(\gamma^{*i},\varepsilon)=\gamma^{*i}(\varepsilon)$ and solves against $\gamma^{*-i}(\varepsilon)$ the optimization problem (7). By combining the above discussion with Lemma 3, we conclude that for any $j\in M_{i},a\in A(I^{j}_{i})$ where $\varpi^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{*}(\varepsilon)$, it holds that $\gamma^{*i}(\varpi^{i}_{I^{j}_{i}}a)=0$. Therefore, $\gamma^{*i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)=y(\gamma^{*i},\varepsilon;\varpi^{i}_{I^{j}_{i}}a)=\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)$. This completes the proof. ∎
@@
Assume there exists a sequence of perturbed games $\{\Gamma_{s}(\varepsilon^{k})\}_{k=1}^{\infty}$ with $\gamma^{*}(\varepsilon^{k})$ being a Nash equilibrium for each $\Gamma_{s}(\varepsilon^{k})$ and $\lim_{k\to\infty}\gamma^{*}(\varepsilon^{k})=\gamma^{*}$. Let $\{\sigma^{k}\}_{k=1}^{\infty}$ be a sequence of totally mixed strategies with $(\sigma^{k},\gamma^{*}(\varepsilon^{k}))\in T$ and $\lim_{k\to\infty}\sigma^{k}=\sigma^{*}$. Note that $\sigma^{k}$ does not necessarily equal $\sigma(\gamma^{*}(\varepsilon^{k}))$. For any pure strategy $s^{i}$ of player $i\in N$, if $u^{i}(s^{i},\sigma^{k-i})<u^{i}(\tilde{s}^{i},\sigma^{k-i})$ holds for some $\tilde{s}^{i}\in S^{i}$, then by Lemma 3, there exists a sequence $\varpi^{i}_{I^{j}_{i}}a$ for $j\in M_{i},a\in A(I^{j}_{i})$ such that $s^{i}(\varpi^{i}_{I^{j}_{i}}a)=1$ and $\varpi^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{*}(\varepsilon^{k})$. Accordingly, Lemma 4 implies that $\gamma^{*i}(\varepsilon^{k};\varpi^{i}_{I^{j}_{i}}a)=\sum_{s^{i}\in S^{i}}s^{i}(\varpi^{i}_{I^{j}_{i}}a)\sigma^{ki}(s^{i})=\eta^{i}(\varepsilon^{k};\varpi^{i}_{I^{j}_{i}}a)$, which leads to $\sigma^{ki}(s^{i})\leq\eta^{i}(\varepsilon^{k};\varpi^{i}_{I^{j}_{i}}a)\leq\varepsilon^{k}$. Therefore the sufficiency follows immediately from Definition 1.
@@
Conversely, assume that $\sigma^{*}$ is a normal-form perfect equilibrium of $\Gamma$ and $\gamma^{*}=\gamma(\sigma^{*})$. Then, there exists a sequence of totally mixed strategies $\{\sigma(\varepsilon^{k})\}_{k=1}^{\infty}$ such that $\lim_{k\to\infty}\varepsilon^{k}=0$ and $\lim_{k\to\infty}\sigma(\varepsilon^{k})=\sigma^{*}$, where each $\sigma(\varepsilon^{k})$ is an $\varepsilon^{k}$-normal form perfect equilibrium. Consider a specific $\sigma(\varepsilon^{k})$ with sufficiently large $k$, if $u^{i}(s^{i},\sigma^{-i}(\varepsilon^{k}))<u^{i}(\tilde{s}^{i},\sigma^{-i}(\varepsilon^{k}))$ holds for some $\tilde{s}^{i}\in S^{i}$, then $\sigma^{i}(\varepsilon^{k};s^{i})\leq\varepsilon^{k}$. Let $\gamma^{k}=\gamma(\sigma(\varepsilon^{k}))$ and $\tilde{\varepsilon}^{k}=\max_{i\in N}|S^{i}|\varepsilon^{k}$, we construct a perturbed game $\Gamma_{s}(\tilde{\varepsilon}^{k})$ with $\gamma^{k}$ being a Nash equilibrium. The perturbation $\eta(\tilde{\varepsilon}^{k})=(\eta^{i}(\tilde{\varepsilon}^{k};\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ adheres to the system (4), defined for $(j,a)\in D_{i}$ as follows,
[…]
@@
It can be observed that $0\leq\eta^{i}(\tilde{\varepsilon}^{k};\varpi^{i}_{I^{j}_{i}}a)\leq\tilde{\varepsilon}^{k}$. Furthermore, the recursive expressions of $\eta(\tilde{\varepsilon}^{k})$ and $\gamma^{k}$ ensure that $\eta^{i}(\tilde{\varepsilon}^{k};\varpi^{i}_{I^{j}_{i}}a)=\gamma^{ki}(\varpi^{i}_{I^{j}_{i}}a)$ holds for all $i\in N,j\in M_{i}$, and $a\in A(I^{j}_{i})$, provided that $\varpi^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma^{k}$. Applying Lemma 4, we conclude that $\gamma^{k}$ is a Nash equilibrium for $\Gamma_{s}(\tilde{\varepsilon}^{k})$. As $\lim_{k\to\infty}\gamma^{k}=\gamma^{*}$, the proof is complete. ∎
@@ 4.1 An Artificial Game and Equilibrium Convergence Analysis
Let $\varepsilon=1/\max_{i\in N,j\in M_{i}}|A(I^{j}_{i})|$ and $\eta^{0}=(\eta^{0i}(\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ be a given vector that satisfies the system (4). Furthermore, let $\gamma^{0}=(\gamma^{0i}(\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ denote a given realization plan profile with $\gamma^{0i}(\varpi^{i})\geq\eta^{0i}(\varpi^{i})$, which serves as a starting point for the smooth path discussed later. The proposed method also requires the following functions, defined over the interval $[0,2]$,
@@
$\begin{array}[]{l}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ 0\leq\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;j\in M_{i},a\in A(I^{j}_{i}),\end{array}$ (8)
@@
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-c(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}(t))\\ &+c(t)\sum\limits_{(j,a)\in D_{i}}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a))\\ &+\theta(t)\sum\limits_{(j,a)\notin D_{i}}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)\ln(\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (9)
@@
$\begin{array}[]{l}(1-c(t))g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(t))+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}\\ \hskip 85.35826pt+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ (\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a))\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=c(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a),\\ \hskip 28.45274pt\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a)<\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;0<\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i},\\ \gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a),\\ \hskip 91.04872pt0<\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;0\leq\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\notin D_{i},\end{array}$ (10)
where $\zeta^{i}_{I^{j}_{i}}(a)=\sum_{{j_{q}}\in M_{i}(\varpi^{i}_{I^{j}_{i}}a)}\nu^{i}_{I^{j_{q}}_{i}}$. One can note that $\gamma^{*}(t)$ constitutes a solution to the optimization problem (9) against itself precisely when there exists a unique pair $(\lambda^{*},\nu^{*})$, along with $\gamma^{*}(t)$, satisfies the system (10). For values of $t\in(0,1]$, the condition $\gamma^{*}(t)\in\text{int}(\Lambda)$ ensures that $\sigma(\gamma^{*}(t))$ is well-defined and $(\sigma(\gamma^{*}(t)),\gamma^{*}(t))\in T$.
@@
Let $V$ denote the set of $v\in\mathbb{R}^{n_{0}}$ satisfying $f_{1}(v)\leq 0,\cdots,f_{l}(v)\leq 0,p_{1}(v)=0,\cdots,p_{q}(v)=0$, where each $f_{i}$ and $p_{j}$ is a polynomial with real coefficients. Suppose that $V$ is nonempty. Then there exist constants $\tau>0,\kappa>0$ and $\kappa^{\prime}>0$ such that $dist(v,V)\leq\tau(1+\|v\|)^{\kappa^{\prime}}(\|[f(v)]_{+}\|+\|p(v)\|)^{\kappa}$ for any $v\in\mathbb{R}^{n_{0}}$. Here $dist(\cdot,\cdot)$ denotes the Euclidean distance function between two sets, $f(v)=(f_{1}(v),\cdots,f_{l}(v))^{\top}$, $p(v)=(p_{1}(v),\cdots,p_{q}(v))^{\top}$, and $\cdot$$_{+}$ denotes the positive part of a vector.